### v3

@@ 4 Wireless Network Large Model
Large-scale AI models (LAM) exhibit remarkable capabilities and great potential in academia and industries. However, such a trend presents new challenges for the underlying network infrastructure. Wireless networks play a vital role in the infrastructure and are directly associated with LAM’s performance, such as data transmission, processing speed, etc. Therefore, this chapter aims to explore the critical role of wireless networks in supporting the operation of these LAM and conduct a systematic analysis and reflection on the relevant issues related to wireless networks in the context of constructing LAM.
@@ 4.1 Comparisons Between LLM and Wireless Network Large Model
LLMs[227] have demonstrated their promise in text and image tasks. It is challenging to directly apply traditional LLMs to wireless communication systems, as the data in the system consists of various protocol data in access networks, CNs, and application data from the upper layer. Furthermore, these data are both structured and non-structured and may be heterogeneous with temporal and relational features.
@@
Specifically, the air interface involves wireless channels’ transmission and scheduling management, encompassing diverse information generated by various wireless communication standards, such as LTE and 5G. The CN handles routing, authentication, and service control of user data, which includes sensitive content such as user behavior and location information, demanding high levels of privacy and security. Operational management data covers status monitoring, fault diagnosis, and performance optimization of network equipment. These datasets are typically large-scale, high-dimensional, lacking effective designs for unifying data across different standards. These characteristics necessitate the consideration of complex factors such as spatiotemporal relationships, real-time requirements, and system stability when processing and analyzing wireless communication data, presenting significant differences and challenges compared to conventional LLM data processing approaches, as illustrated in Figure 22.
@@
Figure 22: Differences between LLM and wireless network large model
Moreover, wireless communication systems constitute a highly complex ecosystem[228] that employs multiple technological paradigms to meet diverse technical requirements and environmental scenarios. For example, coding and modulation techniques are pivotal in ensuring data transmission reliability and efficiency; MIMO technology enhances data transfer rates and signal quality by utilizing multiple transmit and receive antennas; wireless interference management aims to increase overall channel capacity; traffic forecasting facilitates effective network planning and resource allocation; load balancing techniques optimize the distribution of user and data traffic across different network resources; and internet protocol (IP) route optimization maximizes the efficiency of data packet transmission paths. These disparate wireless task models exhibit significant differences, and their commonalities are not always clear, contributing to the need for a nuanced approach to developing a large wireless network model that diverges from conventional LLM methodologies.
@@
In contrast to the relatively lenient requirements for inference speed in LLM models used in text and image processing, the demands placed on model inference speed in wireless communication tasks are critically important. Insufficient inference speed in a large wireless network model can lead to outdated results due to significant changes in channel conditions and signal characteristics during the inference process. Therefore, ensuring that models in wireless communication systems can quickly obtain inference results is a KPI for deployment effectiveness.
@@
On the other hand, the requirements for inference accuracy in large wireless network models far exceed those of LLM. In language models, different combinations of word arrangements have minimal impact on semantic expression; however, in wireless communication systems, which rely on exact mathematical models, sensitivity to inference errors is markedly high, with repercussions that are difficult to quantify. Signal modulation, a critical process encoding information onto carrier signals, exemplifies this sensitivity. Minor deviations in predicting modulation parameters by the model, such as errors in frequency or phase, can prevent the receiver from correctly demodulating signals, resulting in data loss or errors. In distributed wireless networks, even slight errors in predicting clock offsets or sync signals by the model can lead to synchronization failures among nodes, thereby compromising network stability. Therefore, precise control over the inference accuracy of the wireless network large model is a critical research challenge.
@@
Although LLM models have significantly advanced text and image processing, directly applying them to construct a large wireless network model is impractical. In practical applications, integrating domain expertise with advanced data processing techniques is essential for effectively handling and analyzing wireless communication data, addressing its complexities and dynamic challenges.
@@ 4.2 Potential Approaches to Constructing a Wireless Network Large Model
Wireless network large model[229] refers to an intelligent and efficient solution for communication network operations, achieved through integrating knowledge and technologies from the communication domain. The wireless network large model provides network management and operations fault detection and diagnostic services. The failures in networks may lead to performance degradation and service interruptions. By analyzing massive network data, the wireless network large model can monitor network status in a real-time fashion, quickly identify and locate potential reasons, and provide accurate diagnosis strategies.
@@
Furthermore, the wireless network large model optimizes task orchestration and scheduling. Modern communication networks involve complex and diverse tasks and business processes, requiring efficient scheduling and management to ensure network operations. Through an in-depth analysis of network topology and business requirements, the wireless network large model optimizes task execution sequences and resource allocations, enhancing network resource utilization and business processing efficiency.
@@
In terms of performance optimization and resource allocation [230], the wireless network large model employs intelligent strategies and algorithms. By continuously monitoring network load and traffic conditions and integrating predictive analytics and dynamic adjustment mechanisms [231], it achieves real-time optimization of network performance and rational allocation of resources, ensuring network stability and reliability. These functionalities enable the Wireless Network Large Model to enhance network management efficiency and capabilities for operators and network administrators, delivering stable and high-quality communication services to users.
@@
The prevailing industry viewpoint suggests that the wireless network large model should adopt a three-layer model structure, as demonstrated in Figure 23. The L2 layer can provide customized solutions tailored to specific business scenarios in this framework. For instance, the L2 layer model can optimize energy utilization efficiency in energy conservation within networks by analyzing data traffic and device loads. Regarding traffic prediction and fault monitoring, the L2 layer model can employ DL techniques to identify anomalies and issue early warnings, thereby enhancing network stability and reliability.
@@
Subsequently, the L1 layer involves domain-specific grand models. The design of the L1 layer emphasizes modeling and optimization for specific domain challenges and issues. For example, the L1 layer model can analyze wireless channel conditions and user behavior in the air interface to improve data transmission rates and coverage. The L1 layer model can utilize network topology and device status information for fault diagnosis and optimization scheduling in the CN and operations. Lastly, the L0 layer represents the universal grand model for the entire network. The design of this layer needs to consider the overall architecture and operational mechanisms of communication systems to ensure it possesses strong generalization capabilities across various scenarios. Through comprehensive data analysis and intelligent decision-making, it aims to provide optimization and management services for all communication tasks.
Figure 23: The three-layer architecture of a wireless network large model
@@
As shown in Figure 24, the conceptual framework for constructing a large wireless network model in communication networks encompasses architecture design, dataset construction, computational support, and application scenarios. These aspects complement each other to form the fundamental framework of the Wireless Network Large Model.
Figure 24: Potential Approaches
Firstly, considering the architecture design of the Wireless Network Large Model in communication networks, constructing a universally applicable large model for the entire network faces numerous challenges. On the one hand, data within communication networks exhibits highly structured characteristics, encompassing extensive information from various devices and sensors[232], including network traffic, device status, and user behavior. The diversity and complexity of these data present significant challenges for model construction and optimization. On the other hand, tasks within communication networks exhibit complex interdependencies and relationships, with difficulties in clearly defining their commonalities and differences. For instance, network traffic characteristics may vary substantially across different business scenarios, rendering conventional unified modeling approaches less suitable. Therefore, directly constructing a universally applicable large-scale model for the entire network is exceptionally challenging.
@@
A phased approach to model construction becomes particularly crucial to address the aforementioned challenges effectively. This strategy reduces the complexity of model design and optimization and enhances scalability and adaptability, providing reliable foundational support for developing and optimizing large-scale network models. When implementing the phased exploration strategy, the initial focus should be small-scale, offline L2-level models. Through the in-depth research and practical validation of critical business scenarios such as network energy efficiency, traffic prediction, and fault monitoring, efficient utilization of network resources can be achieved, thereby improving overall network performance and sustainability. Subsequently, progression towards large-scale, real-time responsive L1 and L0-level models should be undertaken. In designing L1-level models, emphasis should be placed on addressing domain-specific challenges and issues, employing targeted modeling and optimization approaches to achieve domain generalization capabilities across air interface, operations, and CN sides. To construct a universally applicable L0-level model encompassing the entire network, carefully considering the communication system’s overall architecture and operational mechanisms is essential to ensure the model’s ability to support services and strategies for all communication tasks. Constructing datasets is also crucial for large-scale network models[233] directly impacting AI models’ training efficacy and application performance. Current challenges primarily revolve around data acquisition and quality difficulties [234]. Addressing these issues requires effective measures to ensure dataset quality and completeness.
To tackle the challenge of data acquisition difficulties, multiple avenues should be explored, including proactive collaboration with industry partners to promote data sharing and openness[235], involving telecommunications operators, equipment manufacturers, and others to expand the coverage of datasets. Furthermore, establishing industry standards and guidelines can guide stakeholders in adhering to unified data collection and processing standards, facilitating data sharing and exchange to overcome data acquisition challenges.
In terms of data quality, it is essential to establish a comprehensive data quality assessment framework[236] to conduct thorough evaluations and inspections of the data. Utilizing data quality metrics such as accuracy, completeness, and consistency, the data should be assessed and monitored to detect any data quality anomalies and take corresponding measures for improvement and optimization. Additionally, employing data quality management tools and technologies can automate quality checks and validations, thereby enhancing the efficiency and precision of data quality management.
During the data collection process, it is crucial to enhance the monitoring and management of data sources to ensure the comprehensiveness and reliability of data acquisition. Advanced data cleaning and preprocessing techniques should be implemented to denoise and handle missing values in raw data, enhancing data accuracy and usability. Simultaneously, establishing robust data standardization procedures to unify data formats reduces dataset heterogeneity and inconsistencies, thereby improving the effectiveness of model training.
@@
Regarding the computational demands of 6G communication to support large-scale network models, optimization efforts are needed in both AI training and deployment. This aims to achieve efficient, reliable, secure, and stable data transmission and utilization of computing resources. In AI training services[237], key factors include efficient and reliable training data collection and transmission processes, proprietary protocol design, support from network elements, and customized traffic scheduling algorithms. Proprietary protocols tailored explicitly for AI task transmission should be developed in the design of access and CNs in the design of access and CNs. These protocols need to account for the unique characteristics of AI tasks, such as large data volumes and computational intensity, to ensure optimal transmission efficiency and performance. Secondly, critical network elements must fully support AI task transmission, encompassing network and computing devices. These devices should have sufficient computing and storage capabilities to handle the extensive data and complex computational demands generated during AI training.
When designing network elements to support AI task transmission, considerations should be made regarding hardware architecture and software functionality to process and transmit AI training tasks effectively. Additionally, in traffic scheduling, selecting appropriate strategies is crucial for supporting the training of large AI models. Conventional traffic scheduling strategies may not suffice for AI training tasks, necessitating the design of customized traffic scheduling algorithms tailored to the characteristics of AI tasks. These algorithms should factor in task priorities, network load conditions, and device resource availability to ensure that AI training tasks receive sufficient network and computing resources promptly, thereby facilitating smooth training processes.
@@
In AI inference services, the high demands of large AI models on storage space and computational capabilities often exceed what a single BS can provide. A viable approach is to consider designing an effective architecture utilizing distributed node collaboration [238] to collectively handle model storage and inference tasks. Alongside the design of distributed node collaboration architecture, effective management and scheduling of inference tasks are crucial to maximize the utilization of each node’s computational resources while ensuring system efficiency and stability [239]. A key consideration in this regard is the design of task allocation and scheduling algorithms.
conventional task scheduling algorithms may fail to adequately account for load conditions and computational resource utilization among nodes, resulting in decreased system performance or uneven workload distribution. Hence, intelligent task scheduling algorithms must be designed to dynamically monitor and analyze load conditions across nodes and evaluate inference task priorities and computational resource requirements in real time. This approach allows for the rational allocation and scheduling of inference tasks, effectively enhancing overall system performance and ensuring optimal utilization of node resources to achieve efficient completion of inference tasks.
Additionally, attention should be given to enhancing system monitoring and automated management. Automated management tools and mechanisms, such as automated configuration and deployment, fault diagnosis and recovery, can be introduced to detect and promptly address faults and anomalies. This reduces management and maintenance costs and improves system maintainability and manageability. These optimization measures collectively provide better technical support and assurance for AI large model computation capabilities in 6G communication networks.
4.3 Challenges of Deploying Wireless Network Large Model
@@
Data Heterogeneity: In wireless networks, the diversity and complexity of data present significant challenges in constructing adequate large-scale training and testing corpus. There exists a wide range of structured and unstructured [240] data in networks. These data include uplink and downlink channel information, measurement reports, Minimization of Drive Test data, KPI monitoring data, network topology performance metrics, as well as cost and energy consumption logs. They are highly heterogeneous and exhibit significant differences in granularity and temporal resolution from diverse sources. A primary research concern is effectively aligning such heterogeneous data to establish a consistent and representative dataset. Exploring techniques such as data normalization, time-series alignment, and data fusion can help to build high-quality training corpus. Furthermore, for constructing a large wireless network model, attention must be given to converting processed data into suitable input formats for inference algorithms. This may involve feature extraction and selection, data dimensionality reduction, and constructing appropriate input representations to enable efficient inference and decision-making by the model.
@@
Real-Time Inference: Wireless network large model can be employed for real-time decision-making tasks[241] such as dynamic spectrum allocation, power control, and user access management. However, the long inference time of these large models [242] may degrade overall network performance and user experience. This is particularly critical in network resource scheduling, where stringent time constraints necessitate model inference completion within milliseconds to effectively manage and allocate communication network resources. This directly influences the network’s real-time responsiveness and service quality. To meet these requirements[243], the design and deployment of the wireless network large model must prioritize optimizing inference speed. This involves hardware acceleration, algorithm optimization, and adjustments to network architecture[244]. Moreover, achieving a balance between model accuracy and computational efficiency is essential to enable efficient and reliable management and scheduling of network resources.
@@
Ambiguous Route: It is challenging to build a generalized large-model framework for wireless communications. At the same time, current industry practices favor leveraging large-scale pre-trained language models as foundational elements, subsequently fine-tuned for specific industry applications. The complexity and dynamics of wireless networks demand models with heightened adaptability and flexibility. Recent explorations in semantic communication-based models[245] have made strides towards a more profound understanding and processing of data flows within wireless networks. However, the feasibility and efficacy of these models in practical deployment, including their stability and performance across diverse environments and conditions, remain open research questions requiring further empirical validation.
@@
Protocol Compatibility: Integrating and deploying the wireless network large model inevitably relies on the protocol stack of communication networks. However, since large-model technologies have emerged recently, compatibility issues with existing communication protocol stacks require urgent resolution. The capability of current communication protocol stack architectures to meet the demands of future large-scale models and strategies for augmenting and optimizing existing protocol stacks to support these models in wireless networks better constitute significant technical challenges. This involves thoroughly analyzing existing protocol stacks to identify potential barriers to integration with large-scale models and exploring new protocol design principles to ensure efficient and stable operation of Wireless Network Large Models in wireless network environments.
@@
Security Threats: Wireless network large model faces multiple security challenges [246]. Firstly, these models may suffer from privacy attacks. Datasets collected from 5G networks may involve sensitive information such as user locations, mobility patterns, and communication content. Without adequate protection, these data face privacy threats [247], leading to compromised user trust and legal issues. Secondly, data tampering presents another significant threat to large models. Attackers may inject misleading data or manipulate sources to influence the model’s learning process and decision outcomes. For instance, incorrect input data could misjudge user demands or network congestion conditions. Additionally, network attacks targeting large wireless network models are hazardous. Malicious actors may disrupt model operations or regular requests through flooding or resource consumption attacks, severely impacting service availability and real-time decision-making capabilities. Finally, ensuring the security and integrity of large model architectures is crucial. Reverse engineering and leakage of model parameters could provide attackers with sensitive information about intellectual property or operational mechanisms, jeopardizing service stability and reliability.

### v4

@@ 4 Wireless Network Large Model
LAMs exhibit remarkable capabilities and great potential in academia and industries. However, such a trend presents new challenges for the underlying network infrastructure. Wireless networks play a vital role in the infrastructure and are directly associated with LAMs’ performance (e.g., data transmission, processing speed, etc). Therefore, this section aims to explore the critical role of wireless networks in supporting the operation of these LAMs and conduct a systematic analysis and reflection on the relevant issues related to wireless networks in the context of constructing LAMs.
@@ 4.1 Comparisons Between LLM and Wireless Network Large Model
LLMs[252] have demonstrated their promise in text and image tasks. It is challenging to directly apply traditional LLMs to wireless communication systems, as the data in the system consists of various protocol data in access networks, CN, and application data from the upper layer. Furthermore, these data are either structured or non-structured and may be heterogeneous with temporal and relational features. Specifically, the air interface involves wireless channels’ transmission and scheduling management, encompassing diverse information generated by various wireless communication standards, such as LTE and 5G. The CN handles routing, authentication, and service control of user data, which includes sensitive content such as user behavior and location information, demanding high levels of privacy and security. Operational management data covers status monitoring, fault diagnosis, and performance optimization of network equipment. These datasets are typically large-scale, high-dimensional, lacking effective designs for unifying data across different standards. These characteristics necessitate the consideration of complex factors such as spatiotemporal relationships, real-time requirements, and system stability when processing and analyzing wireless communication data, presenting significant differences and challenges compared to conventional LLM data processing approaches, as illustrated in Figure 20.
@@
Figure 20: Differences between LLMs and wireless network large models
@@
Moreover, wireless communication systems constitute a highly complex ecosystem[253] that employs multiple technological paradigms to meet diverse technical requirements and environmental scenarios. Unlike the relatively lenient requirements for inference speed in LLMs used in text and image processing, the demands placed on model inference speed in wireless communication tasks are critically important. Insufficient inference speed in a wireless network large model can lead to outdated results due to significant changes in channel conditions and signal characteristics during the inference process. Therefore, ensuring that models in wireless communication systems can quickly obtain inference results is a KPI for deployment effectiveness.
@@
On the other hand, the requirements for inference accuracy in wireless network large models far exceed those of LLMs. In language models, different combinations of word arrangements have minimal impact on semantic expression; however, in wireless communication systems, which rely on exact mathematical models, sensitivity to inference errors is markedly high, with repercussions that are difficult to quantify. Signal modulation, a critical process encoding information onto carrier signals, exemplifies this sensitivity. Minor deviations in predicting modulation parameters by the model, such as errors in frequency or phase, can prevent the receiver from correctly demodulating signals, resulting in data loss or errors. In distributed wireless networks, even slight errors in predicting clock offsets or sync signals by the model can lead to synchronization failures among nodes, thereby compromising network stability. Therefore, precise control over the inference accuracy of the wireless network large model is a critical research challenge.
@@
Although LLMs have significantly advanced text and image processing, directly applying them to construct a wireless network large model is impractical. In practical applications, integrating domain expertise with advanced data processing techniques is essential for effectively handling and analyzing wireless communication data, addressing its complexities and dynamic challenges.
@@ 4.2 Potential Approaches to Constructing a Wireless Network Large Model
Wireless network large model[254] refers to an intelligent and efficient solution to communication network operations, achieved through integrating knowledge and technologies from the communication domain. A wireless network large model provides network management, and operations fault detection and diagnostic services. The failures in networks may lead to performance degradation and service interruptions. By analyzing massive network data, a wireless network large model can monitor network status in a real-time fashion, quickly identify and locate potential reasons, and provide accurate diagnosis strategies.
@@
A wireless network large model optimizes task orchestration and scheduling. Modern communication networks involve complex and diverse tasks and business processes, requiring efficient scheduling and management to ensure network operations. Through an in-depth analysis of network topology and business requirements, a wireless network large model can potentially optimize task execution sequences and resource allocations, enhancing network resource utilization and business processing efficiency.
@@
In terms of performance optimization and resource allocation [255], the wireless network large model employs intelligent strategies and algorithms. By continuously monitoring network load and traffic conditions and integrating predictive analytics and dynamic adjustment mechanisms [256], it achieves real-time optimization of network performance and rational allocation of resources, ensuring network stability and reliability. The prevailing industry viewpoint suggests that the wireless network large model should adopt a three-layer model structure, involving L0, L1, and L2 layers.
@@
  • The L2 layer can provide customized solutions tailored to specific business scenarios in this framework. For instance, the L2 layer model can optimize energy utilization efficiency in energy conservation within networks by analyzing data traffic and device loads. Regarding traffic prediction and fault monitoring, the L2 layer model can employ DL techniques to identify anomalies and issue early warnings, thereby enhancing network stability and reliability.
@@
  • The L1 layer involves domain-specific grand models. The design of the L1 layer emphasizes modeling and optimization for specific domain challenges and issues. For example, the L1 layer model can analyze wireless channel conditions and user behavior in the air interface to improve data transmission rates and coverage. The L1 layer model can utilize network topology and device status information for fault diagnosis and optimization scheduling in the CN and operations.
  • The L0 layer represents the universal grand model for the entire network. The design of this layer needs to consider the overall architecture and operational mechanisms of communication systems to ensure it possesses strong generalization capabilities. Through comprehensive data analysis and intelligent decision-making, it aims to provide optimization and management services for all communication tasks.
@@
Figure 21 shows the construction philosophy for constructing a wireless network large model in communication networks, encompassing evolution route, dataset construction, and computational support, which complement each other to form the fundamental framework of wireless network large models.
Figure 21: Construction philosophy of wireless network large model
@@
Key 1: Evolution Route. Constructing a universally applicable large model for the entire network faces many challenges. On the one hand, data within communication networks exhibit highly structured characteristics, encompassing extensive information from various devices and sensors [257], including network traffic, device status, and user behavior. The diversity and complexity of these data present significant challenges for model optimization. On the other hand, tasks within communication networks can be complex and interdependent. For instance, network traffic characteristics may vary substantially across different business scenarios, rendering conventional unified modeling approaches less feasible. A phased approach to model construction becomes important to address the aforementioned challenges effectively. The initial focus can be small-scale, offline L2-level models. Through the practical validation of critical business scenarios such as network energy efficiency, traffic prediction, and fault monitoring, efficient utilization of network resources can be achieved. In designing L1-level models, emphasis should be placed on addressing domain-specific challenges, employing targeted modeling and optimization approaches to achieve domain generalization capabilities across air interface, operations, and CN sides. To construct a universally applicable L0-level model encompassing the entire network, carefully considering the communication system’s overall architecture and operational mechanisms is essential to ensure the model’s ability to support services and strategies for all communication tasks.
Key 2: Dataset Construction. Dataset Construction is crucial for large-scale network models[258] directly impacting AI model training efficacy and application. Challenges revolve around data acquisition and quality difficulties [259]. To tackle the difficulties in data acquisition, multiple avenues should be explored, including proactive collaboration with industry partners to promote data sharing and openness[260], involving telecommunications operators, equipment manufacturers, and others to expand the coverage of datasets. Furthermore, establishing industry standards and guidelines can guide stakeholders in adhering to unified data collection and processing standards, facilitating data sharing and exchange to overcome data acquisition challenges. A comprehensive data quality assessment framework is also indispensable [261]. Utilizing data quality metrics such as accuracy, completeness, and consistency, data should be assessed and monitored to detect any data quality anomalies and take corresponding measures for improvement and optimization. Employing data quality management tools and technologies can automate quality checks and validations. During the data collection process, it is crucial to enhance the monitoring and management of data sources to ensure the comprehensiveness and reliability of data acquisition. Advanced data cleaning and preprocessing techniques should be implemented to denoise and handle missing values in raw data. Moreover, establishing robust data standardization procedures to unify data formats reduces dataset heterogeneity and inconsistencies.
@@
Key 3: Computational Support. Regarding the computational demands of 6G communication to support large-scale network models, optimization efforts are needed in both AI training and deployment. This aims to achieve efficient, reliable, secure, and stable data transmission and utilization of computing resources. In AI training services, key factors include efficient and reliable training data collection and transmission processes, proprietary protocol design, support from network elements, and customized traffic scheduling algorithms[262]. When designing core and access networks, proprietary protocols tailored specifically for AI task transmission should be developed. These protocols need to account for the unique characteristics of AI tasks, such as large data volumes and computational intensity, to ensure optimal transmission efficiency and performance. When designing network elements to support AI task transmission, considerations should be made regarding hardware architecture and software functionality to process and transmit AI training tasks effectively. In traffic scheduling, selecting appropriate strategies is crucial for supporting the training of LAMs. Conventional traffic scheduling strategies may not suffice for AI training tasks, necessitating the design of customized traffic scheduling algorithms tailored to the characteristics of AI tasks. These algorithms should factor in task priorities, network load conditions, and device resource availability to ensure that AI training tasks receive sufficient network and computing resources promptly.
@@
In AI inference services, the high demands of LAMs on storage space and computational capabilities often exceed what a single BS can provide. A viable approach is designing an effective architecture utilizing distributed node collaboration [263] to collectively handle model storage and inference tasks. Alongside the design of distributed node collaboration architecture, effective management and scheduling of inference tasks are crucial to maximize the utilization of each node’s computational resources while ensuring system efficiency and stability [264]. A key consideration is the design of task allocation and scheduling algorithms. This approach allows for the rational allocation and scheduling of inference tasks. Moreover, attention should be given to enhancing system monitoring and automated management. Automated management tools and mechanisms, such as automated configuration and deployment, fault diagnosis, and recovery, can be introduced to detect and promptly address faults and anomalies. This reduces management and maintenance costs and improves system maintainability and manageability. These optimization measures collectively provide better technical support and assurance for AI large model computation capabilities in 6G networks.
4.3 Challenges of Wireless Network Large Model Development
@@
Data Heterogeneity. In wireless networks, the diversity and complexity of data present significant challenges in constructing comprehensive large-scale training and testing corpora. The network contains various types of structured and unstructured data, including uplink and downlink channel information, measurement reports, drive test data minimization, KPI monitoring data, network topology performance indicators, and cost and energy consumption logs. These data sources are highly heterogeneous, and the data from different origins often exhibit notable discrepancies in granularity and time resolution. It is necessary to consider how to effectively align these heterogeneous data to build a consistent and representative dataset. Exploring techniques such as data normalization [265], time-series alignment [266], and data fusion [267] can contribute to constructing a high-quality training corpus. Additionally, attention should be given to transforming the processed data into input formats suitable for inference algorithms. This may involve feature extraction and selection, data dimensionality reduction, and the creation of appropriate input representations to enable the model to perform efficient inference and decision-making.
@@
Real-Time Inference. Wireless network large model can be applied to real-time decision-making tasks such as dynamic spectrum allocation, power control, and user access management. However, excessive inference time can degrade overall network performance and user experience. This is particularly critical in network resource scheduling, where strict time constraints demand that model inference be completed within milliseconds to effectively manage and allocate communication network resources. This directly impacts the real-time responsiveness and service quality of the network. To meet these requirements, the design and deployment of large-scale models in wireless networks must prioritize optimizing inference speed, which includes hardware acceleration, algorithm optimization, and network architecture adjustments[268]. Furthermore, achieving a balance between model accuracy and computational efficiency is crucial for efficient and reliable network resource management and scheduling .
@@
Ambiguous Route. Building a universal large-scale model framework for wireless communication is a challenging task. Current industry practices tend to leverage large-scale pre-trained language models as foundational components, which are then fine-tuned for specific industry applications. However, the complexity and dynamics of wireless networks require models with higher adaptability and flexibility. Scholars have made some progress in exploring models based on semantic communication, which has advanced the understanding and processing of data flows in wireless networks [269]. However, the feasibility and effectiveness of these models in practical deployments, including their stability and performance under varying environments and conditions, remain unresolved and require further empirical validation.
@@
Protocol Compatibility. The integration and deployment of the Wireless Network Large Model inevitably rely on the communication network’s protocol stack. However, since large-scale model technologies have only emerged in recent years, compatibility issues with existing communication protocol stacks remain to be addressed. Whether the current protocol stack architecture can meet the future demands of large-scale models, and how to enhance and optimize the existing protocol stack to better support the application of these models in wireless networks, pose significant technical challenges. This requires an analysis of the existing protocol stack to identify potential barriers to large-scale model integration and the exploration of new protocol design principles to ensure the efficient and stable operation of large-scale models in wireless networks.
@@
Security Threats. Wireless network large model faces multiple security challenges[270]. Firstly, it may encounter privacy attacks. Data sets from 6G networks may involve sensitive information, such as user location, mobility patterns, and communication content. Without adequate protection, this data is vulnerable to privacy threats, which can lead to a loss of user trust and legal issues. Secondly, data tampering is another significant threat to large-scale models. Attackers may influence the model’s learning process and decision-making outcomes by injecting misleading data or manipulating data sources. For instance, incorrect input data could lead to misjudgments of user demands or network congestion conditions. Additionally, network attacks targeting large-scale wireless network models are also hazardous. Malicious actors may disrupt model operations or regular requests through flooding attacks or resource depletion attacks, severely affecting service availability and real-time decision-making capabilities. Finally, ensuring the security and integrity of large-scale model architectures is critical. Reverse engineering and model parameter leakage could allow attackers to gain sensitive information about intellectual property or operational mechanisms, thereby jeopardizing the stability and reliability of services.